OpenStudy (elisaneedshelp):

Three consecutive odd integers have a sum of 21. Find the integers. :( pleaseee

OpenStudy (aihberkhan):

Okay. he meaning of consecutive means continuous {please don't be offended } so let the first term be 'y', the second one be 'y+1' and the third one be 'y+2' so according to the question, y+(y+1)+(y+2)=21 you don't need to worry as there is a '+' sign outside the bracket you just need to add = y+y+1+y+2=21 group the like terms [terms with same ending like 'y'] = y+y+y+1+2 =21 3y+3=21 {just to tell you y*y*y= y³ } now 3y=21-3 3y=18 y=18/3 y=6 so the integers are y=6, y+1=6+1=7 and y+2= 6+2=8 if you want to check=6+7=13+8=21 hence proved

OpenStudy (aihberkhan):

Hope this helped! Have a great day! Also a medal would be much appreciated! Just click best response next to my answer. Thank You! @ElisaNeedsHelp

OpenStudy (elisaneedshelp):

Ofc the medal :D lol